Suppose that A is a 3x3 matrix and that b is a 3x1 matrix. Assuming that
the matrix equation Ax = b has infinitely many solutions, write down all the possible
reduced row-echelon forms of the augmented matrix (A | b). You may use ‘*’ to
represent an unknown number.
1
Expert's answer
2020-01-30T13:01:33-0500

Consider the matrix

"A=\begin{pmatrix}
 a_{11} &a_{12}&a_{13} \\
 a_{21} &a_{22}&a_{23} \\
a_{31} &a_{32}&a_{33} \\
\end{pmatrix}"

"X=\begin{pmatrix}
 x_1 \\
 x_2\\
x_3
\end{pmatrix}"

"b=\begin{pmatrix}
 b_1 \\
 b_2\\
b_3
\end{pmatrix}"

Subject to the problem the matrix equation

"Ax = b"

has infinitely many solutions then "det A=0, rank(A)<3".

The matrix equation "Ax = b" :


"\left\{\begin{matrix}
 a_{11}x_1+ a_{12}x_2+a_{13}x_3=b_1\\
 a_{21}x_1+ a_{22}x_2+a_{23}x_3=b_2\\
a_{31}x_1+ a_{32}x_2+a_{33}x_3=b_3\\
\end{matrix}\right."

Consider the matrix

"(A|b)=\begin{pmatrix}
 a_{11} &a_{12}&a_{13} &|b_1\\
 a_{21} &a_{22}&a_{23} &|b_2\\
a_{31} &a_{32}&a_{33} &|b_3\\
\end{pmatrix}"

Suppose "a_{11}\neq 0, rank(A)=2"


"(A|b) \leftrightarrow\begin{pmatrix}
 1 &\frac{a_{12}}{a_{11}}&\frac{a_{13}}{a_{11}} &|\frac{b_1}{a_{11}}\\
 a_{21} &a_{22}&a_{23} &|b_2\\
a_{31} &a_{32}&a_{33} &|b_3\\
\end{pmatrix} \leftrightarrow\\
\begin{matrix}
 R_2+R_1(-a_{21}) \\
 R_3+R_1(-a_{31})
\end{matrix}\\
\leftrightarrow
\begin{pmatrix}
 1 &*&* &|*\\
 0 &c_{22}&c_{23} &|c_2\\
0 &c_{32}&c_{33} &|c_3\\
\end{pmatrix}"

"c_{22}\neq 0, rg(A)=2"

perform operations over rows

"\frac{R_2}{c_{22}}\\
R_3+R_2(-c_{32})"


"(A|b)\leftrightarrow \begin{pmatrix}
 1 &*&* &|*\\
 0 &1&* &|*\\
0 &0&0 &|0\\
\end{pmatrix}"

Similarly, if "rank(A)=1, a_{11}\neq 0", then

"(A|b)\leftrightarrow \begin{pmatrix}
 1 &*&* &|*\\
 0 &0&0 &|0\\
0 &0&0 &|0\\
\end{pmatrix}"